OpenStudy (anonymous):

*Solve the differential equation dy/dx=x/16y. *Find the equation of the solution through the point (x,y)=(−4,1). *Find the equation of the solution through the point (x,y)=(0,−4). Your answer should be of the form y=f(x)

OpenStudy (anonymous):

I already solved the first part. C=8y^2 -(x^2/2)

myininaya (myininaya):

ok replace your x and y with -4 and 1 respectively to find the C so that your graph for your equation will go through that point

OpenStudy (anonymous):

so y=sqrt((x^2/16) + c) and i plugged in (-4,1) and got y=sqrt((x^2/16) is that correct?

OpenStudy (anonymous):

c= 0 for the second question?

myininaya (myininaya):

You have \[C=8y^2-\frac{x^2}{2}\] Replace x with -4 and y with 1 like this: \[C=8(1)^2-\frac{(-4)^2}{2}\] Is this what you did to solve for C?

myininaya (myininaya):

\[C=8-\frac{16}{2}=8-8=0\]

OpenStudy (anonymous):

yes that's what i did

myininaya (myininaya):

I can't do arithmetic somethings lol But you have \[0=8y^2-\frac{x^2}{2} \]

myininaya (myininaya):

that is the equation satisfying the point (-4,1)

OpenStudy (anonymous):

hmm it the system doesn't seem to accept it.. it says it isn't a linear equation

OpenStudy (anonymous):

doesn't that mean i have to solve for y=...

myininaya (myininaya):

Your differential equation was: \[\frac{dy}{dx}=\frac{x}{16y} \text{ \right ?} \]

OpenStudy (anonymous):

yup

myininaya (myininaya):

Ok solve for y make sure you keep in mind we want to go through the point (-4,1)

OpenStudy (anonymous):

ok so y=x/4

OpenStudy (anonymous):

but it still doesn't work??

OpenStudy (anonymous):

oh wait y=-x/4 works!!

myininaya (myininaya):

We wanted y=-x/4 That is why I said to keep in mine we want it going through (-4,1) :)

myininaya (myininaya):

mind*

myininaya (myininaya):

The equation y=x/4 doesn't satisfy the point (-4,1) but y=-x/4 does

myininaya (myininaya):

Like 1=-(-4)/4

OpenStudy (anonymous):

ohh i get it now

OpenStudy (anonymous):

thanks a lot!!

OpenStudy (anonymous):

i'm going to do the last part now

OpenStudy (anonymous):

c=128 for the last part right?

OpenStudy (anonymous):

oh i got the last part too!!

OpenStudy (anonymous):

thanks a lot again!!!

myininaya (myininaya):

yes C=128 for the last one